# FERGI: Automatic Annotation of User Preferences for Text-to-Image Generation from Spontaneous Facial Expression Reaction

### 05 Dec 2023 v1

@@ Abstract
Researchers have proposed to use data of human preference feedback to fine-tune text-to-image generative models. However, the scalability of human feedback collection has been limited by its reliance on manual annotation. Therefore, we develop and test a method to automatically annotate user preferences from their spontaneous facial expression reaction to the generated images. We collect a dataset of Facial Expression Reaction to Generated Images (FERGI) and show that the activations of multiple facial action units (AUs) are highly correlated with user evaluations of the generated images. Specifically, AU4 (brow lowerer) is most consistently reflective of negative evaluations of the generated image. This can be useful in two ways. Firstly, we can automatically annotate user preferences between image pairs with substantial difference in AU4 responses to them with an accuracy significantly outperforming state-of-the-art scoring models. Secondly, directly integrating the AU4 responses with the scoring models improves their consistency with human preferences. Additionally, the AU4 response best reflects the user’s evaluation of the image fidelity, making it complementary to the state-of-the-art scoring models, which are generally better at reflecting image-text alignment. Finally, this method of automatic annotation with facial expression analysis can be potentially generalized to other generation tasks. The code is available at https://github.com/ShuangquanFeng/FERGI, and the dataset is also available at the same link for research purposes.
@@ 1 Introduction
We present the Facial Expression Reaction to Generated Images (FERGI) dataset, which comprises video recordings of 33 participants’ facial expression reaction to 2827 images generated by Stable Diffusion v1.4 [58] based on 576 different self-drafted text prompts, along with their feedback on the generated images from manual input. We estimate the activation of facial action units (AUs), as defined in the Facial Action Coding System (FACS, a comprehensive system breaking down facial expressions into individual components of AUs of muscle movements [19]), in the facial expression reaction videos with a pre-trained model and show that the activations of multiple AUs are highly correlated with user evaluations of the generated images. We further propose a method of automatically annotating user preferences by detecting the activation of AU4 in the facial expression reaction to the generated images and show that this evaluation metric best reflects the user’s evaluation of the image fidelity, making it complementary to the state-of-the-art scoring models, which are generally better at reflecting image-text alignment.
@@ 2.2 Data Collection Procedure
Each participant completes multiple sessions of data collection. In each session, the participant drafts one text prompt and views 5 images generated from the input text prompt. A flow chart of the procedure of each session is shown in Fig. 1. Details are explained below.
@@ 2.2.2 Recording
After the input text prompt is finalized, the participant is directed to a webcam preview to confirm that the webcam captures their face appropriately. The webcam is then used to record throughout the rest of the session. The configuration of the raw videos vary based on the participant’s device and browser. Though, recorded videos of most participants have a resolution of $640\times 480$, and all recorded videos are standardized to 30 FPS afterwards.
@@ 2.2.3 Image Presentation, Annotation, and Ranking
In each session, a total of 5 images generated by Stable Diffusion v1.4 [58] from the input text prompt are presented and annotated sequentially and ranked afterwards (present image 1 $\rightarrow$ annotate image 1 $\rightarrow$ $\dots$ $\rightarrow$ present image 5 $\rightarrow$ annotate image 5 $\rightarrow$ rank 5 images). The image presentation and annotation follows the webcam preview. The detailed procedures for the presentation and annotation of each generated image are as follows:
@@
  • Annotate (no time restriction): The website displays the prompt, the generated image (moved to the right side), and the annotation survey. The participant annotates the image manually by filling out the survey during this interval. The survey was adapted from the survey designed for ImageReward [73] and includes the following questions: important elements not reflected in the image (free-response question), overall rating (on a scale of 1 (worst) to 7 (best)), image-text alignment rating (on a scale of 1 to 7), fidelity rating (on a scale of 1 to 7), issues of the image [^1] For the first 7 out of 33 participants (or 8 out of the total 39 participants), there was a bug (which was fixed afterwards) in checkbox selection for answering the questions regarding “issues of the image” that made the recorded answers unreliable, so the data regarding their answers for this question should not be used in analysis. (multi-choice question), and emotions felt when seeing the image (multi-choice question).
After the annotation of the last image, the participant ranks the 5 generated images from best to worst, and the data of the session is uploaded to the web server.
@@ 2.3 Data Preprocessing
Stable Diffusion v1.4 [58] outputs an all-black image when it detects the original output image might be inappropriate. A total of 53 such images out of 2880 total generated were generated in the dataset. The data associated with these images were excluded from further analysis, resulting in less than 5 valid images for some input text prompts.
@@ 3 AU Model Training
We analyze the facial expression reaction of the users by estimating the activation of their AUs. We trained our AU estimation model on the DISFA [49, 48] and DISFA+ [47] datasets. In these datasets, each frame is manually annotated by a human expert with intensities on a scale of 0 to 5, for activation of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop). Video frames are first preprocessed with face detection [44], face alignment [6], and a combination of histogram equalization and linear mapping [35] before being fed into the model for training, testing, or inference. For model training, we use the neural network IR-50 [11] pre-trained on Glink360k [1] and fine-tune it on the DISFA and DISFA+ datasets [49, 47]. A single network learns the estimation of all AUs labeled in the datasets with both regression and ordinal classification [51].
@@ 4 Facial Feature Extraction
In this section, we elaborate the procedure of extracting facial features from the video clips associated with each generated image based on the trained AU recognition model. Similar procedures were used to extract features from baseline clips and reaction clips.
@@ 4.1 Data Filtering
Robustness against face occlusion and faces with non-frontal poses is an ongoing challenge for facial expression recognition models [79, 27]. Our AU model was trained on datasets consisting predominantly of frontal facial images without occlusion and has not been specifically adapted for non-frontal faces or faces with occlusion, so it is not suitable for estimating AUs of such faces. We address this issue by excluding such data from analysis. Specifically, for each 5-second video clip, we exclude frames with a low face detection confidence score (given by MediaPipe [44]) and frames with an off-angle pitch or yaw (estimated based on detected facial landmarks [6]). Then, if a clip has more than $20\%$ of its total frames being excluded, the whole clip, and its associated data (survey and ranking results) are excluded from further analysis.
@@ 4.2 AU Activation Value
We compute an AU activation value $\alpha_{i}$ for each trained AU (with $i$ representing the index of the AU) in each video clip. The AU activation value is computed as follows:
@@ 5.1 Statistical Analysis
  • Overall Ratings. The participant’s overall ratings (on a scale of 1 to 7) of the generated images have a significant negative correlation with the activation values of AU4 and AU9 as shown in Figs. 1(a) and 1(b) (Spearman correlation: $\rho\approx-0.18$ and $p<5\times 10^{-17}$ for AU4, $\rho\approx-0.09$ and $p<5\times 10^{-5}$ for AU9).
@@
  • Emotions. For the question “Did you feel any of the following emotions when you saw the image?”, the participant can decide whether to include the six options (disappointed, satisfied, surprised, disgusted, amused, scared) in the answer independently (they can select 0 to 6 of the options as their answer). We found that whether they reportedly feel each of the emotions are significantly correlated with the activation values of multiple AUs.
    • Disappointed. ”Disappointing” images are associated with lower activation values of AU1, AU2, and AU12 and higher activation values of AU4 as shown in Fig. 2(a) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
    • Satisfied. ”Satisfying” images are associated with lower activation values of AU4, AU5, AU9, and AU25 as shown in Fig. 2(b) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).
    • Surprised. ”Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Fig. 2(c) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
    • Disgusted. ”Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Fig. 2(d) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
    • Amused. ”Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Fig. 2(e) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
    • Scared. ”Scary” images are associated with higher activation values of AU4 and AU5 as shown in Fig. 2(f) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).
@@
5.2 Automatic Annotation of User Preferences between Image Pairs Using AU4
We have observed that AU4 activation value has the highest and most consistent (negative) correlation with the raw overall ratings in the previous subsection. Therefore, we use the participants’ AU4 activation values to automatically annotate their preferences between images generated based on the same input text prompt.
@@ 5.2.1 Participant Exclusion
Multiple factors may result in systematically unreliable, unstable AU4 estimation from a person by our AU model. We exclude 3 participants with unreliable AU4 estimation from the analysis of automatic preference annotation with the following criterion: a participant is excluded if the 80th percentile of the AU4 activation values in their baseline clips exceeds 0.5. (Actually, the lowest among the excluded ones is 0.74 while the highest among the included ones is 0.36.) Among the three excluded participants, two of them have bangs covering the area between the eyebrows while the other one has highly unstable camera angles in the videos.
5.2.2 AU4 Valence Score
The AU4 activation value $\alpha_{4}$ is a good indicator of the user’s evaluation of the generated image, but the raw value might not perfectly reflect the associated evaluation. Therefore, we propose an AU4 valence score $s_{\mathrm{AU4}}$, designed to have a more linear relationship and positive correlation with the associated evaluation of the generated image. We define the AU4 valence score $s_{\mathrm{AU4}}$ as the negative of an exponential decay function of AU4 activation value $\alpha_{4}$ as follows:
$s_{\mathrm{AU4}}=-(1-e^{-k\cdot\alpha_{4}}),$ (3)
where $k$ is an exponential decay coefficient to fit.

### 21 Mar 2024 v2

@@ Abstract
Researchers have proposed to use data of human preference feedback to fine-tune text-to-image generative models. However, the scalability of human feedback collection has been limited by its reliance on manual annotation. Therefore, we develop and test a method to automatically annotate user preferences from their spontaneous facial expression reaction to the generated images. We collect a dataset of Facial Expression Reaction to Generated Images (FERGI) and show that the activations of multiple facial action units (AUs) are highly correlated with user evaluations of the generated images. Specifically, AU4 (brow lowerer) is reflective of negative evaluations of the generated image whereas AU12 (lip corner puller) is reflective of positive evaluations. These can be useful in two ways. Firstly, we can automatically annotate user preferences between image pairs with substantial difference in these AU responses with an accuracy significantly outperforming state-of-the-art scoring models. Secondly, directly integrating the AU responses with the scoring models improves their consistency with human preferences. Finally, this method of automatic annotation with facial expression analysis can be potentially generalized to other generation tasks. The code is available at https://github.com/ShuangquanFeng/FERGI, and the dataset is also available at the same link for research purposes.
@@ 1 Introduction
We present the Facial Expression Reaction to Generated Images (FERGI) dataset, which comprises video recordings of 33 participants’ facial expression reaction to 2827 images generated by Stable Diffusion (SD) v1.4 [58] based on 576 different self-drafted text prompts, along with their feedback on the generated images from manual input. We estimate the activation of facial action units (AUs), as defined in the Facial Action Coding System (FACS, a comprehensive system breaking down facial expressions into individual components of AUs of muscle movements [20]), in the facial expression reaction videos with a pre-trained model and show that the activations of multiple AUs are highly correlated with user evaluations of the generated images. We further propose a method of automatically annotating user preferences by detecting the activations of AU4 (brow lowerer, see Fig. 5) and AU12 (lip corner puller, see Fig. 6) in the facial expression reaction to the generated images.
Figure 1: A flow chart summarizing our proposed pipeline. The solid lines represent the procedures demonstrated in the paper while the dashed lines represent the procedures proposed but not demonstrated. Specifically, we developed a model and method for computing an AUcomb valence score and integrating it with state-of-the-art text-to-image scoring models. We propose that the AUcomb valence score could be used to fine-tune/improve these text-to-image scoring models and that the integrated ensemble score could be used to fine-tune the text-to-image generative models.
To better illustrate the idea of our paper, we present Fig. 1 showing a flow chart summarizing the pipeline the paper proposed.
2 Related Work
Facial Expression Recognition and Applications. Automatic facial expression recognition has advanced rapidly in recent years [42]. As emotional responses to image generation can be more complicated than single categories of the most widely-researched seven basic emotions [19, 18, 47], we decided to analyze the facial expression reactions directly in terms of muscle movements, as defined by AUs in FACS [20]. AU detection and estimation has attracted increasing interest [27, 61, 84, 83, 11, 43, 44, 69, 34, 35, 85, 46, 65, 66] and been used for facial emotion recognition [77, 79, 78, 76]. Specifically related to our usage of AU4 and AU12 to infer user evaluations of generated images, researchers have shown that the activation of the corrugator (activated in AU4) and zygomaticus (activated in AU12) are associated with negative and positive emotions respectively and are the most investigated muscle activations in studying affect and emotion [72]. Also, the activity of the corrugator supercilii (activated in AU4) is positively associated with amygdala and negatively associated with ventromedial prefrontal cortex activity [24] and increases with negative stimuli, such as negative images [6, 5, 38, 37]. For applications, [3, 71] showed the effectiveness of using facial expression to infer/analyze human preferences.
@@ 3.2 Data Collection Procedure
Each participant completed multiple sessions of data collection. In each session, the participant drafted one text prompt and viewed 5 images generated from the input text prompt. A flow chart of the procedure of each session is shown in Fig. 2. Details are explained below.
@@ 3.2.2 Recording
After the input text prompt is finalized, the participant is directed to a webcam preview to confirm that the webcam captures their face appropriately. The webcam is then used to record throughout the rest of the session. The configuration of the raw videos varied based on the participant’s device and browser. Though, recorded videos of most participants have a resolution of $640\times 480$, and all recorded videos were standardized to 30 FPS afterwards.
@@ 3.2.3 Image Presentation, Annotation, and Ranking
In each session, a total of 5 images generated by SD v1.4 [58] from the input text prompt are presented and annotated sequentially and ranked afterwards (present image 1 $\rightarrow$ annotate image 1 $\rightarrow$ $\dots$ $\rightarrow$ present image 5 $\rightarrow$ annotate image 5 $\rightarrow$ rank 5 images). The image presentation and annotation followed the webcam preview. The detailed procedures for the presentation and annotation of each generated image are as follows:
@@
  • Annotate (no time restriction): The website displays the prompt, the generated image (moved to the right side), and the annotation survey. The participant annotates the image manually by filling out the survey during this interval. The survey was adapted from the survey designed for ImageReward [75] and includes the following questions: important elements not reflected in the image (free-response question), overall rating (on a scale of 1 (worst) to 7 (best)), image-text alignment rating (on a scale of 1 to 7), fidelity rating (on a scale of 1 to 7), issues of the image [^1] For the first 7 out of 33 participants (or 8 out of the total 39 participants), there was a bug (later fixed) in checkbox selection for answering the questions regarding “issues of the image” that made the recorded answers unreliable, so the data regarding their answers for this question should not be used in analysis. (multi-choice question), and emotions felt when seeing the image (multi-choice question).
After the annotation of the last image, the participant ranks the 5 generated images from best to worst, and the data of the session is uploaded to the web server. That marks the end of the session.
@@ 3.3 Data Preprocessing
SD v1.4 [58] outputs an all-black image when it detects the original output image might be inappropriate. A total of 53 such images out of 2880 total generated were generated in the dataset. The data associated with these images were excluded from further analysis, resulting in fewer than 5 valid images for some input text prompts.
@@ 4 AU Model Training
We analyze the facial expression reaction of the users by estimating the activation of their AUs. We trained our AU estimation model on the DISFA [50, 49] and DISFA+ [48] datasets. They are among the largest, most popular, and most cited AU datasets with intensity annotations. In these datasets, each frame is manually annotated by a human expert with intensities on a scale of 0 to 5, for activation of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop). Video frames are first preprocessed with face detection [45], face alignment [7], and a combination of histogram equalization and linear mapping [36] before being fed into the model for training, testing, or inference. For model training, we used the neural network IR-50 [12] pre-trained on Glink360k [1] and fine-tuned it on the DISFA and DISFA+ datasets [50, 48]. A single network learned the estimation of all labeled AUs with both regression and ordinal classification [52]. More details of the training of the AU model is elaborated in the Supplementary Materials.
@@ 5 Facial Feature Extraction
In this section, we elaborate the procedure of extracting facial features from the video clips associated with each generated image based on the trained AU recognition model. Similar procedures were used to extract features from baseline clips and reaction clips. Further details are in the Supplemental Materials.
@@ 5.1 Data Filtering
Robustness against face occlusion and faces with non-frontal poses is an ongoing challenge for facial expression recognition models [81, 28]. Our AU model was trained on datasets consisting predominantly of frontal facial images without occlusion and thus for each 5-second video clip, we exclude frames with a low face detection confidence score (given by MediaPipe [45]) and frames with an off-angle pitch or yaw (estimated based on detected facial landmarks [7]). Then, if a clip has more than $20\%$ of its total frames being excluded, the whole clip, and its associated data (survey and ranking results) are excluded from further analysis.
@@ 5.2 AU Activation Value
We compute an AU activation value $\alpha_{i}$ for each trained AUi (with $i$ representing the index of the AU) in each video clip. The AU activation value is computed as follows:
@@ 6.1 Statistical Analysis
  • Overall Ratings. The participant’s overall ratings (on a scale of 1 to 7) of the generated images have a significant positive correlation with the activation values of AU2 and AU12 and a significant negative correlation with the activation values of AU4 and AU9 as shown in Figs. 2(f), 2(a), 2(b) and 2(g) (Spearman correlation: $\rho\approx 0.09$ and $p<5\times 10^{-5}$ for AU2, $\rho\approx-0.18$ and $p<5\times 10^{-17}$ for AU4, $\rho\approx-0.09$ and $p<5\times 10^{-5}$ for AU9, $\rho\approx 0.08$ and $p<5\times 10^{-4}$ for AU12).
@@
  • Emotions. For the question “Did you feel any of the following emotions when you saw the image?”, the participant can independently select 0 to 6 answers from the six options: disappointed, satisfied, surprised, disgusted, amused, and scared. Each of the reported emotions was found to be significantly correlated with the activation values of multiple AUs.
    • Disappointed. “Disappointing” images are associated with lower activation values of AU1, AU2, and AU12 and higher activation values of AU4 as shown in Fig. 3(a) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
    • Satisfied. “Satisfying” images are associated with lower activation values of AU4, AU5, AU9, and AU25 as shown in Fig. 3(b) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).
    • Surprised. “Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Fig. 3(c) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
    • Disgusted. “Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Fig. 3(d) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
    • Amused. “Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Fig. 3(e) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
    • Scared. “Scary” images are associated with higher activation values of AU4 and AU5 as shown in Fig. 3(f) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).
@@
6.2 Automatic Annotation of User Preferences between Image Pairs Using AU4 and AU12
Figure 5: An example of AU4 activation in response to a low-quality image generation. (a) The input text prompt. (b) The facial expression of the participant before seeing the generated image. (c) The generated image. (d) The facial expression of the participant after seeing the generated image (the frame with the highest estimated AU4 intensity). In the annotation survey, the participant gave an overall rating of 2, an image-text alignment rating of 2, and a fidelity rating of 1; she identified the image with the issues “output contains unwanted content that was not mentioned in the text prompt” and “existence of body problem”; she reported that she felt “disappointed” when seeing the image; she ranked the image as the worst among the five images generated from the same prompt.
Figure 6: An example of AU12 activation in response to a high-quality image generation. (a) The input text prompt. (b) The facial expression of the participant before seeing the generated image. (c) The generated image. (d) The facial expression of the participant after seeing the generated image (the frame with the highest estimated AU12 intensity). In the annotation survey, the participant gave ratings of 7 for each of image-text alignment, fidelity, and overall rating; he did not identify any issues with the image; he reported feeling “satisfied”, “surprised”, and “amused” when seeing the image; he ranked the image as the best among the five images generated from the same prompt.
We aim to use the estimated AU activation values to automatically annotate user preferences between images generated from the same input text prompt. We have observed that the activations of AU2, AU4, AU9, and AU12 are significantly correlated with the the raw overall ratings in the previous subsection. After comparing combinations between them, we developed the following pipeline of using AU4 and AU12 to do automatic annotation (examples of AU4 and AU12 responses are shown in Figs. 5 and 6). (Inclusion of AU2 and AU9 did not help.) We also did not find a way to effectively incorporate the information conveyed by AUs that correlate only with the extremity of the ratings. Collection of a much larger-scale dataset in the future might enable effective incorporation of such information with a more complex model.
@@ 6.2.1 Participant Exclusion
While estimations of AU12 are generally reliable and stable, multiple factors may result in unreliable AU4 estimation from a small number of people by our AU model. We excluded 3 participants with unreliable AU4 estimation due to bangs covering the area between the eyebrows (2 participants) or highly unstable camera angles (1 participant).
6.2.2 AUcomb Valence Score
The AU4 activation value $\alpha_{4}$ and AU12 activation value $\alpha_{12}$ are good indicators of the user’s evaluation of the generated image, but the raw values might not linearly reflect the associated evaluation. Therefore, we propose an AU4 valence score $s_{\mathrm{AU4}}$ and an AU12 valence score $s_{\mathrm{AU12}}$, designed to have a more linear relationship and positive correlation with the associated evaluation of the generated image. We define the $s_{\mathrm{AU4}}$ as the negative of an exponential decay function of AU4 activation value $\alpha_{4}$ and $s_{\mathrm{AU12}}$ as the positive exponential decay function of AU12 activation value $\alpha_{12}$ as follows:
$s_{\mathrm{AU4}}=-(1-e^{-k\cdot\alpha_{4}})$ (3)
and
$s_{\mathrm{AU12}}=(1-e^{-k\cdot\alpha_{12}}),$ (4)
where $k$ is an exponential decay coefficient to fit. We further define the final AUcomb valence score $s_{\mathrm{AUcomb}}$ as the linear combination of them:
$s_{\mathrm{AUcomb}}=w_{\mathrm{AU4}}s_{\mathrm{AU4}}+w_{\mathrm{AU12}}s_{\mathrm{AU12}}$ (5)
subject $\text{subject to}\quad w_{\mathrm{AU4}}+w_{\mathrm{AU12}}$ $=1,$ (6)
where $w_{\mathrm{AU4}}$ and $w_{\mathrm{AU12}}$ represent the weights for AU4 and AU12 respectively.